The equation of the directrix of the parabola y2+4y+4x+2=0 is:

Grade:12

1 Answers

Vijay Mukati
8 years ago
Dear Student,

Given equation of parabola can be rewritten as (y+2)2 = -4(x-0.5).
Now compairing which Y2 = 4aX,
a = -1.
Therefore, the equation of directric is given by X + a =0
Therfore (x – 0.5) + (-1) = 0
x – 1.5 = 0.

Thanks.